OpenStudy (anonymous):

The list of a certain tool is x dollars. In store A the original selling price of the tool was $50 less than the list price, and the current selling price is 10% less than the original selling price. In store B the original selling price of the tool was 10% less than the list price, and the current selling price is $50 less than the original selling price. A) Quality A is greater B) Quality B is greater C) The two quantities are equal

OpenStudy (anonymous):

(x-50) then 10% of that = your answer (x-10%)-50= your answer

OpenStudy (anonymous):

ok so say that x=100 A: 100-50= 50 then take 10% 50 therefore it =5 B: 100-10%-50 i think it would be 10-50 which would =-40. ok now I' confused. Sorry. just try to set up the B equation and plug in a # for x and you will find your answer

OpenStudy (anonymous):

thanks!